### v1

@@ Abstract
Context. Observations have revealed that the Milky Way, Andromeda, Centaurus A (and potentially other galaxies) host spatially thin and kinematically coherent planes of satellites. Such structures are highly improbable within the standard $\Lambda$CDM cosmological model, and the dynamical stability of these planes has been a subject of debate for a long time. Accurately determining their stability requires a thorough understanding of orbital parameters such as proper motion, distance, and line-of-sight velocity, in addition to the gravitational potential of the host galaxy. However, many of these remain insufficiently constrained, leading to significant uncertainties in any analysis.
Aims. This research aims to explore the impact of measurement errors in proper motions and distances of the satellite galaxies and in the adopted host halo mass on the inferred stability of satellite planes in Milky-Way-like potentials.
Methods. Test satellite galaxies orbiting a host galaxy are simulated, mock observed by adding various degrees and types of observational errors, and then backward-integrated. Trends and correlations between the initial conditions and the applied uncertainties on the inferred orbital stability of the satellite systems are analyzed. Additionally, the effects of adopting incorrect potentials and the impact of different orbital eccentricities are considered.
Results. Uncertainties in proper motions lead to an inferred, ostensible widening of an intrinsically stable satellite plane, with its width increasing linearly with the adopted proper motion uncertainties. Even uncertainties on the level of Gaia systematics strongly affect the plane’s inferred past width. Moreover, the potential with a low halo mass showed a significant impact on the stability of these planes, while the remaining two host models showed similar effects. Uncertainties in satellite distance also contribute noticeably to the inferred, apparent instability.
@@ 1 Introduction
A century ago, astronomers were acquainted with only two prominent satellite galaxies orbiting the Milky Way, the Large Magellanic Cloud (LMC) and the Small Magellanic Cloud (SMC) – a knowledge that had persisted for centuries (von Humboldt & Otte, 1997). This celestial census doubled when Shapley (1938) utilized photographic plates and discovered two more dwarfs – Sculptor and Fornax. With the discovery of Sagittarius in 1994, the number of satellite galaxies of the Milky Way reached 11 (Ibata et al., 1994), and has substantially increased due to advancements in observational technology, the deployment of larger telescopes, and systematic sky surveys. Currently, approximately 60 dwarf galaxies have been discovered around the Milky Way (Simon, 2019). Notably, it has also been discovered that these bright satellite galaxies of the Milky Way tend to lie on a great circle called the “Magellanic Plane” (Lynden-Bell, 1976). Observations have shown that satellite galaxies around the Milky Way are highly anisotropic and kinematically correlated, with most of them being located in a thin plane roughly perpendicular to the disc of the Milky Way (Kroupa et al., 2005; Metz et al., 2008). The most luminous among these galaxies appear to align within a thin disk-like structure known as the Vast Polar Structure (VPOS, Pawlowski et al. 2012; Taibi et al. 2024). The VPOS exhibits a root mean square (rms) thickness of 20-30 kpc, a radius of about 250 kpc and an axis ratio $c/a$ of 0.18 – 0.30 (Pawlowski, 2018, 2021b).
@@
A comparable thin and anisotropic distribution of satellite galaxies around Andromeda, referred to as The Great Plane of Andromeda (GPoA) (Ibata et al., 2013; Conn et al., 2013), and Centaurus A, known as Centaurus A Satellite Plane (CASP) (Tully et al., 2015; Müller et al., 2018; Kanehisa et al., 2023b), was also found. These structures are characterized by an axis ratio of $c/a$ = 0.1 (Pawlowski et al., 2013) and $c/a$ ratio = 0.2 (Tully et al., 2015), respectively. Both planes, GPoA and CASP, are reported to have a significant correlation in the line-of-sight velocities of satellite members when observed along the edge-on orientation (Ibata et al., 2013; Müller et al., 2021). This suggests that satellites exhibit a consistent sense of co-rotation, with satellites on one side of the host predominantly receding and those on the other side approaching relative to the host. The rotating pattern around its host galaxy, akin to the VPOS, thus supports the idea that both the GPoA and CASP constitute a nonrandom and coherent structure. Moreover, the first proper motion measurements for three on-plane satellite galaxies of M31 indicate that these are indeed consistent with co-orbiting along the spatially identified GPoA (Sohn et al., 2020; Pawlowski & Tony Sohn, 2021; Casetti-Dinescu et al., 2024). A similar flattened structure with coherent line-of-sight velocities has recently been discovered around the galaxy NGC 4490 and found to be rare in cosmological simulations (Karachentsev & Kroupa, 2024; Pawlowski et al., 2024).
1.1 Plane Setup
@@
Figure 1: Face-on (left) and Edge-on (right) views of $N_{\text{sat}}$ = 25 randomly generated test satellite points represented by different colors. The star symbol represents the host galaxy.
@@
Today, the Lambda Cold Dark Matter ($\Lambda$CDM) cosmological model is widely accepted among researchers due to its ability to explain a range of astrophysical phenomena such as Big Bang Nucleosynthesis, the accelerated expansion of the universe (Riess et al., 1998), the power spectrum of the Cosmic Microwave Background (Spergel et al., 2003), large-scale structure (Peacock et al., 2001; Tegmark et al., 2004) and the matter-energy budget of our universe (Ade et al., 2016). However, on smaller scales it presents several challenges. Among them, the Plane of Satellite Galaxies has persistently been a topic of debate (Pawlowski, 2018) and it has been found that such structures are considered highly improbable within the standard $\Lambda$CDM cosmological model (Kroupa et al., 2005; Pawlowski, 2021b). Several solutions have been suggested to explain this discrepancy; however, none have achieved general consensus. These proposed solutions include the accretion along filaments of the cosmic web (Libeskind et al., 2011; Lovell et al., 2011), in-fall of satellite galaxies in groups (Li & Helmi, 2008; D’Onghia & Lake, 2008; Metz et al., 2009; Shao et al., 2018; Wang et al., 2013; Júlio et al., 2024), galaxy mergers (Smith et al., 2016; Kanehisa et al., 2023a), effects induced by the infall of a massive LMC (Garavito-Camargo et al., 2021; Pawlowski et al., 2022; Vasiliev, 2023), tidal dwarf galaxies (Pawlowski et al., 2011; Fouquet et al., 2012; Banik et al., 2022), and hydrodynamics and baryonic physics (Pawlowski et al., 2015; Ahmed et al., 2017; Müller et al., 2018; Pawlowski et al., 2019). Thus, while each hypothesis contributes valuable insights, a unified explanation remains elusive, highlighting the complexity of galaxy formation and evolution processes.
@@
Investigation of the orbits of satellite galaxies with precision requires a thorough understanding of several critical parameters. These parameters encompass precise measurements of the satellite’s position, distance, line-of-sight velocity, proper motions, the underlying gravitational potential of the host galaxy, and the Galactocentric distance of the Sun, along with its motion relative to the Galactic center. Unfortunately, many of these key parameters remain inadequately constrained, introducing a notable degree of uncertainty in orbital analyses (Bland-Hawthorn & Gerhard, 2016; Joshi, 2007; Schönrich et al., 2010). An essential consideration is measurement uncertainties, especially in proper motions, which can significantly affect analyses (Pawlowski, 2021a). These uncertainties can be broadly classified into statistical and systematic uncertainties (Li et al., 2021). Statistical uncertainties, arising from random measurement variations, can be mitigated through statistical methods and the utilization of larger data sets. Conversely, systematic uncertainties, not stemming from random variations but rather from inherent biases in the measurement process or instruments, pose more fundamental challenges.
@@
Maji et al. (2017) demonstrate that the Milky Way’s disk of satellites (DoS) is dynamically transient. Their orbit modeling shows that the 11 classical satellites move away from the current DoS, causing it to thicken from $c/a\sim 0.18$ (height $\sim 19.6$ kpc) to $c/a\sim 0.36$ (45 kpc) in 0.5 Gyr and $c/a\sim 0.42$ (64 kpc) in 1 Gyr. This evolution persists across the two tested galactic potentials:one replacing the stellar disk with a single-component Miyamoto–Nagai potential and another using only an NFW dark matter halo potential. However, Pawlowski et al. (2017) notes that Maji et al. (2017) have overlooked measurement uncertainties, such as those related to proper motions in their studies, and argues that ignoring observational biases further undermines the reliability of Maji et al. (2017)’s conclusions regarding the plane of satellite galaxies. Similarly, Sawala et al. (2023) argues that the plane of satellites is more likely transient rather than rotationally supported, because a back- and forward-integration of satellite galaxy orbits resulted in a widening of the inferred plane thickness. However, they did not take the systematic errors in Gaia proper motions into account. These uncertainties are crucial, as they can significantly impact the measured values and potentially alter the results. Without accounting for these uncertainties, the conclusions drawn about the stability and alignment of the Milky Way’s satellite plane might be overly optimistic or pessimistic. This could lead to misinterpretations about the dynamical status of these satellite structures, thereby influencing our understanding of the formation and evolution of the Milky Way and its satellite system. The incorporation of measurement uncertainties, particularly in proper motion data, is essential for achieving a more realistic and reliable understanding of the satellite galaxies’ orbital behaviors. This is especially the case when studying the joint behavior of an ensemble of satellite galaxies, since errors then have the effect of washing out intrinsic correlations (and approaches such as Monte-Carlo sampling from the uncertainties effectively apply the errors twice, Pawlowski 2021a).
@@
To address this, in this paper, we explore how different levels and types of measurement uncertainties affect the inferred stability of planes of satellite galaxies. Specifically, we employ computer simulations to model the orbital evolution of test satellite galaxies around a host galaxy that were set up as intrinsically stable satellite planes. We mock-observed these by applying controlled degrees of observational errors on their evolved positions and velocities, and then backward-integrate them to determine how well the intrinsically stable dynamical evolution can be inferred.
@@
This paper is structured as follows. In Sect. 1 we explain the methodology. This includes setting up an artificial satellite plane, the employed plane fitting routine, the performed orbit integrations, how mock measurement errors were applied, and the subsequent backward integration of the test satellites. Our results are compiled in Sect. 3 for various combinations of measurement errors, assumed Milky Way potentials, and initial orbital eccentricity distributions. We end with a discussion and conclusions in Sect. 4.
@@ 2 Methodology
To study the evolution of artificial planes of satellite galaxies, it is essential that the test satellites are appropriately arranged around the host galaxy to initially form a planar structure. To achieve this, we are using the Milky Way and its satellite distribution as a model to establish an artificial plane of satellite galaxies. The Milky Way’s satellite plane, VPOS, spans approximately 250 kpc in radius with a root mean square thickness ranging from 20 to 30 kpc. Therefore, the radial distribution, $R$, of $N_{\text{sat}}$ test satellites ranges from 20 to 250 kpc, while the vertical distribution extends from 0 to 20 kpc (Pawlowski, 2018, 2021b). To generate a random radial distribution, we utilize the Von Neumann rejection algorithm, also referred to as rejection sampling (Neumann, 1951). This statistical technique facilitates the generation of random samples from a desired probability distribution. Fig. 1 shows the face-on and edge-on views of the randomly distributed $N_{\text{sat}}=25$ test satellites around a host galaxy, which is initially set up as a planar structure. Here, the satellite plane is set to be perpendicular to the host galaxy’s stellar disk. To produce the radial distance distribution, a power-law distribution function is used:
$f(r)=r^{\alpha}$ (1)
Here, for $\alpha$, we used $-3$. For the velocities, first we need the unit vectors for each test satellite in the radial, tangential, and perpendicular directions, with respect to the satellite plane. In vector representation, they can be written as:
@@
where, $\mathbf{V_{\text{rad; i}}}$, $\mathbf{V_{\text{tan; i}}}$, and $\mathbf{V_{\text{perp; i}}}$ represent the radial, tangential, and perpendicular velocity vectors of the $i$-th test satellite, respectively. $|\mathbf{V_{\text{rad; i}}}|$, $|\mathbf{V_{\text{tan; i}}}|$, and $|\mathbf{V_{\text{perp; i}}}|$ denote the magnitudes of the radial, tangential, and perpendicular velocity, respectively. $\hat{r}_{i}$, $\hat{t}_{i}$, and $\hat{p}_{i}$ signify the unit vectors along the radial, tangential, and perpendicular directions for the $i$-th test satellite.
Hammer et al. (2021) and Cautun & Frenk (2017) investigated the relations between radial and tangential velocities of Milky Way satellite galaxies and reported a tangential bias. Specifically, Cautun & Frenk (2017) demonstrated that 9 out of 10 satellites with measured proper motions exhibit highly tangentially biased motions, with 80% or more of their orbital kinetic energy attributed to tangential motion. Given the significance of tangential velocity, we set both $|\mathbf{V_{\text{perp; i}}}|$ and $|\mathbf{V_{\text{rad; i}}}|$ to zero initially, indicating no motion in radial and perpendicular directions. This means only the tangential component is relevant. Satellites are initialized with a tangential speed equal to the $MW_{\text{fiducial}}$, see Table 1, gravitational potential’s circular velocity, guaranteeing circular orbits. To make orbits eccentric, the tangential unit vectors of all satellites are rotated by an angle $\theta$ around the perpendicular unit vector. For each test satellite, the values for $\theta$ are drawn from a uniform distribution with some maximum range of $-\theta_{\mathrm{max}}$ and $+\theta_{\mathrm{max}}$, as mentioned in Table 2.
Once we have the radial $\mathbf{V_{\text{rad; i}}}$, perpendicular $\mathbf{V_{\text{perp; i}}}$, and tangential $\mathbf{V_{\text{tan; i}}}$ velocity components, we convert them into a Cartesian coordinate system. Thus, this allows us to fully describe the motion of test satellites in the simulated galaxy system.

### v2

@@ Abstract
Context. Observations have revealed that the Milky Way, Andromeda, Centaurus A, and possibly other galaxies host spatially thin and kinematically coherent planes of satellites. Such structures are highly improbable within the standard $\Lambda$CDM cosmological model, and the dynamical stability of these planes has long been debated. Accurately determining their stability requires a thorough understanding of orbital parameters such as proper motion, distance, and line-of-sight velocity, in addition to as the gravitational potential of the host galaxy. However, many of these parameters remain poorly constrained, leading to significant uncertainties in analyses.
Aims. This study explores the impact of measurement errors in the proper motions and distances of the satellite galaxies and in the adopted host halo mass on the inferred stability of satellite planes in Milky Way-like potentials.
Methods. We simulated mock-observed test satellite galaxies orbiting a host galaxy by adding various degrees and types of observational errors, and then backward-integrated the orbits. We analyzed trends and correlations between the initial conditions and the uncertainties applied on the inferred orbital stability of the satellite systems. We also considered the effects of adopting incorrect potentials and the impact of different orbital eccentricities.
Results. Uncertainties in proper motions lead to an apparent widening of an intrinsically stable satellite plane, with its width increasing linearly with the uncertainties in the adopted proper motion. Even uncertainties at the level of Gaia systematics strongly affect the plane’s inferred past width. Moreover, the potential with a low halo mass has a significant impact on the stability of these planes, whereas the remaining two host models show similar effects. Uncertainties in satellite distance also contribute noticeably to the inferred instability.
@@ 1 Introduction
A century ago, astronomers were aware of only two prominent satellite galaxies orbiting the Milky Way, the Large Magellanic Cloud (LMC) and the Small Magellanic Cloud (SMC), a view that had persisted for centuries (von Humboldt & Otte, 1997). This census doubled when Shapley (1938) used photographic plates to discover two additional dwarfs, Sculptor and Fornax. With the discovery of Sagittarius in 1994, the number of known Milky Way satellite galaxies reached 11 (Ibata et al., 1994), and has since increased substantially due to advances in observational technology, the deployment of larger telescopes, and systematic sky surveys. Currently, approximately 60 dwarf galaxies have been discovered around the Milky Way (Simon, 2019). Notably, these bright satellite galaxies tend to lie along a great circle known as the “Magellanic Plane” (Lynden-Bell, 1976). Observations show that Milky Way satellite galaxies are highly anisotropic and kinematically correlated, with most located in a thin plane roughly perpendicular to the disk of the Milky Way (Kroupa et al., 2005; Metz et al., 2008). The most luminous among these galaxies appear to align within a thin, disk-like structure known as the Vast Polar Structure (VPOS; Pawlowski et al. 2012; Taibi et al. 2024). The VPOS exhibits a root mean square (rms) thickness of 20-30 kpc, a radius of about 250 kpc, and an axis ratio $c/a$ of 0.18 – 0.30 (Pawlowski, 2018, 2021b).
@@
Similarly, a comparable thin and anisotropic distribution of satellite galaxies is also found around Andromeda, referred to as the Great Plane of Andromeda (GPoA) (Ibata et al., 2013; Conn et al., 2013), and around Centaurus A, known as the Centaurus A Satellite Plane (CASP) (Tully et al., 2015; Müller et al., 2018; Kanehisa et al., 2023b). These structures are characterized by an axis ratio of $c/a$ = 0.1 (Pawlowski et al., 2013) and a $c/a$ ratio = 0.2 (Tully et al., 2015), respectively. Both the GPoA and CASP planes exhibit significant correlations in the line-of-sight velocities of their satellite members when observed edge-on (Ibata et al., 2013; Müller et al., 2021). This suggests that satellites corotate, with satellites on one side of the host predominantly receding and those on the other side approaching relative to the host. This rotating pattern around its host galaxy, similar to the VPOS, supports the idea that both the GPoA and CASP constitute a nonrandom, coherent structures. Moreover, the first proper motion measurements for three on-plane satellite galaxies of M31 indicate that these galaxies are consistent with co-orbiting along the spatially identified GPoA (Sohn et al., 2020; Pawlowski & Tony Sohn, 2021; Casetti-Dinescu et al., 2024). A similar flattened structure with coherent line-of-sight velocities was discovered around the galaxy NGC 4490 and is found to be rare in cosmological simulations (Karachentsev & Kroupa, 2024; Pawlowski et al., 2024).
@@
Figure 1: Face-on (left) and edge-on (right) views of $N_{\text{sat}}$ = 25 randomly generated test satellites, each represented by a different color. The star symbol denotes the host galaxy.
@@
Today, the Lambda cold dark matter ($\Lambda$CDM) cosmological model is widely accepted among researchers due to its ability to explain a range of astrophysical phenomena such as Big Bang nucleosynthesis, the accelerated expansion of the Universe (Riess et al., 1998), the power spectrum of the cosmic microwave background (Spergel et al., 2003), large-scale structure (Peacock et al., 2001; Tegmark et al., 2004), and the matter-energy budget of our Universe (Ade et al., 2016). However, on smaller scales, it presents several challenges. Among these, the plane of satellite galaxies has consistently remained a topic of debate (Pawlowski, 2018), and such structures are considered highly improbable within the standard $\Lambda$CDM cosmological model (Kroupa et al., 2005; Pawlowski, 2021b). Several solutions have been proposed to explain this discrepancy; however, no general consensus has been reached. These proposed solutions include accretion along filaments of the cosmic web (Libeskind et al., 2011; Lovell et al., 2011), infall of satellite galaxies in groups (Li & Helmi, 2008; D’Onghia & Lake, 2008; Metz et al., 2009; Shao et al., 2018; Wang et al., 2013; Júlio et al., 2024), galaxy mergers (Smith et al., 2016; Kanehisa et al., 2023a), effects induced by the infall of a massive LMC (Garavito-Camargo et al., 2021; Pawlowski et al., 2022; Vasiliev, 2023), tidal dwarf galaxies (Pawlowski et al., 2011; Fouquet et al., 2012; Banik et al., 2022), and hydrodynamics and baryonic physics (Pawlowski et al., 2015; Ahmed et al., 2017; Müller et al., 2018; Pawlowski et al., 2019). Thus, while each hypothesis contributes valuable insight, a unified explanation remains elusive, highlighting the complexity of galaxy formation and evolution processes.
@@
Precise investigation of the orbits of satellite galaxies requires a thorough understanding of several critical parameters. These parameters encompass precise measurements of a satellite’s position, distance, line-of-sight velocity, proper motions, the underlying gravitational potential of the host galaxy, the Galactocentric distance of the Sun, and its motion relative to the Galactic center. However, many of these key parameters remain inadequately constrained, introducing a significant uncertainty in orbital analyses (Bland-Hawthorn & Gerhard, 2016; Joshi, 2007; Schönrich et al., 2010). Measurement uncertainties, especially in proper motions, can significantly affect analyses (Pawlowski, 2021a). These uncertainties can be broadly classified as statistical or systematic (Li et al., 2021). Statistical uncertainties arising from random measurement variations can be mitigated through statistical methods and the use of larger datasets. Conversely, systematic uncertainties, which result from inherent biases in the measurement process or instruments, pose more fundamental challenges.
@@
Maji et al. (2017) demonstrate that the Milky Way’s disk of satellites (DoS) is dynamically transient. Their orbit modeling shows that the 11 classical satellites move away from the current DoS, causing it to thicken from $c/a\sim 0.18$ (height $\sim 19.6$ kpc) to $c/a\sim 0.36$ (45 kpc) in 0.5 Gyr, and $c/a\sim 0.42$ (64 kpc) in 1 Gyr. This evolution persists across the two tested galactic potentials: one replacing the stellar disk with a single-component Miyamoto–Nagai potential and another using only an Navarro–Frenk–White (NFW) dark matter halo potential. However, Pawlowski et al. (2017) notes that Maji et al. (2017) overlook measurement uncertainties, such as those related to proper motions, and argues that ignoring observational biases further undermines the reliability of their conclusions regarding the plane of satellite galaxies. Similarly, Sawala et al. (2023) argues that the plane of satellites is more likely to be transient rather than rotationally supported, because backward and forward integration of satellite galaxy orbits results in a widening of the inferred plane thickness. However, they do not account for systematic errors in Gaia proper motions. These uncertainties are crucial because they can significantly affect the measured values and potentially alter the results. Without accounting for these uncertainties, the conclusions drawn about the stability and alignment of the Milky Way’s satellite plane might be overly optimistic or pessimistic. This could lead to misinterpretations about the dynamical status of these satellite structures, thereby influencing our understanding of the formation and evolution of the Milky Way and its satellite system. Incorporating measurement uncertainties, particularly in proper motion data, is essential to achieve a more realistic and reliable understanding of the satellite galaxies’ orbital behaviors. This is especially true when studying the joint behavior of an ensemble of satellite galaxies, since errors can conceal out intrinsic correlations (and approaches such as Monte Carlo sampling from the uncertainties effectively apply the errors twice; Pawlowski 2021a).
@@
To address this, we explore how different levels and types of measurement uncertainties affect the inferred stability of planes of satellite galaxies. Specifically, we employed computer simulations to model the orbital evolution of test satellite galaxies around a host galaxy that were set up as intrinsically stable satellite planes. We mock-observed these by applying controlled degrees of observational errors on their evolved positions and velocities, and then backward-integrated them to determine how well the intrinsically stable dynamical evolution can be inferred.
@@
This paper is structured as follows. In Sect. 1, we describe the methodology, including setting up an artificial satellite plane, the plane-fitting routine, the orbit integrations performed, the application of mock measurement errors, and the subsequent backward integration of the test satellites. Section 3 presents our results for various combinations of measurement errors, assumed Milky Way potentials, and initial orbital eccentricity distributions. We end with a discussion and conclusions in Sect. 4.
@@ 2 Methodology
2.1 Plane setup
To study the evolution of artificial planes of satellite galaxies, the test satellites must be appropriately arranged around the host galaxy to initially form a planar structure. To achieve this, we used the Milky Way and its satellite distribution as a model to establish an artificial plane of satellite galaxies. The Milky Way satellite plane, VPOS, spans approximately 250 kpc in radius, with a root mean square thickness ranging from 20 to 30 kpc. Therefore, the radial distribution, $R$, of $N_{\text{sat}}$ test satellites ranges from 20 to 250 kpc, while the vertical distribution extends from 0 to 20 kpc (Pawlowski, 2018, 2021b). To generate a random radial distribution, we used the Von Neumann rejection algorithm, also referred to as rejection sampling (Neumann, 1951). This statistical technique facilitates the generation of random samples from a desired probability distribution. Figure 1 shows the face-on and edge-on views of the randomly distributed $N_{\text{sat}}=25$ test satellites around a host galaxy, initially set up as a planar structure. Here, the satellite plane is set perpendicular to the host galaxy’s stellar disk. To produce the radial distance distribution, we used a power-law distribution function:
$f(r)=r^{\alpha}.$ (1)
Here, for $\alpha$, we used $-3$. For the velocities, we first defined the unit vectors for each test satellite in the radial, tangential, and perpendicular directions, with respect to the satellite plane. In vector representation, these are written as
@@ 2.1 Plane setup
where $\mathbf{V_{\text{rad; i}}}$, $\mathbf{V_{\text{tan; i}}}$, and $\mathbf{V_{\text{perp; i}}}$ represent the radial, tangential, and perpendicular velocity vectors of the $i$-th test satellite, respectively. The quantities $|\mathbf{V_{\text{rad; i}}}|$, $|\mathbf{V_{\text{tan; i}}}|$, and $|\mathbf{V_{\text{perp; i}}}|$ denote the magnitudes of the radial, tangential, and perpendicular velocity, respectively. The unit vectors $\hat{r}_{i}$, $\hat{t}_{i}$, and $\hat{p}_{i}$ correspond to the radial, tangential, and perpendicular directions for the $i$-th test satellite.
Hammer et al. (2021) and Cautun & Frenk (2017) investigated the relations between radial and tangential velocities of Milky Way satellite galaxies and report a tangential bias. Specifically, Cautun & Frenk (2017) show that nine out of ten satellites with measured proper motions exhibit highly tangentially biased motions, with 80% or more of their orbital kinetic energy attributed to tangential motion. Given the significance of tangential velocity, we initially set both $|\mathbf{V_{\text{perp; i}}}|$ and $|\mathbf{V_{\text{rad; i}}}|$ to zero, indicating no motion in the radial and perpendicular directions. Thus, only the tangential component is relevant. Satellites were initialized with a tangential speed equal to the circular velocity of the $MW_{\text{fiducial}}$ gravitational potential (see Table 1), guaranteeing circular orbits. To make the orbits eccentric, we rotated the tangential unit vectors of all satellites by an angle $\theta$ around the perpendicular unit vector. For each test satellite, the values for $\theta$ were drawn from a uniform distribution with a maximum range of $-\theta_{\mathrm{max}}$ and $+\theta_{\mathrm{max}}$, as listed in Table 2.
Once we obtained the radial $\mathbf{V_{\text{rad; i}}}$, perpendicular $\mathbf{V_{\text{perp; i}}}$, and tangential $\mathbf{V_{\text{tan; i}}}$ velocity components, we converted them into a Cartesian coordinate system. This allowed us to fully describe the motion of the test satellites in the simulated galaxy system.